OpenStudy (anonymous):

An isosceles triangle has area of 110 ft². If the base is 14 ft, what is the length of the legs? Round your answer to the nearest tenth. A 21 ft B 17.2 ft C 14.8 ft D 442.9 ft

OpenStudy (solomonzelman):



OpenStudy (solomonzelman):

the area of this triangle would be: altitude times 14 divided by 2

OpenStudy (solomonzelman):

(I will disregard the units for now) so, you are given the area to be 110: (let altitude =a) \(\large\color{slate}{ 110=a\times14\div2 }\) can you solve for a?

OpenStudy (anonymous):

No I'm so confused right now.

OpenStudy (solomonzelman):

okay, do you get that: \(\LARGE\color{teal}{ A_{\LARGE \rm ~ isosceles~\triangle}=({\LARGE \rm altitude} \cdot {\LARGE \rm ~ base}) \div2 }\) ??

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

now, lets plug in the given information. you know: base=14 area=110

OpenStudy (solomonzelman):

\(\LARGE\color{teal}{ A_{\LARGE \rm ~ isosceles~\triangle}=({\LARGE \rm altitude} \cdot {\LARGE \rm ~ base}) \div2 }\) PLUGGING IN ..... \(\LARGE\color{teal}{ 100=({\LARGE \rm altitude} \cdot {\LARGE \rm ~ 14}) \div2 }\)

OpenStudy (solomonzelman):

can you solve for altitude?

OpenStudy (anonymous):

No, I'm not sure how to get the altitude

OpenStudy (triciaal):